y=6x-11
-2x-3y=-7

Found 2 solutions by richwmiller, TheProdicalSon:
Answer by richwmiller(9144)

-2x-3y=-7
y=6x-11
substitute
-2x-3(6x-11)=-7
-2x-18x+33=-7
combine
-20x=-40
divide both sidesby -20
x=2
y=6(2)-11
y=1
check
-2(2)-3(1)=-7
-4-3=-7
-7=-7
ok

Answer by TheProdicalSon(33)

y = 6x - 11
-2x - 3y = -7
Get everything on the correct side
y - 6x = -11
-3y - 2x = -7
Multiply the top equation by 3 to eliminate the y
3y - 18x = -33
-3y - 2x = -7
Add the equations together
-20x = -40
x = 2
Insert x into one of the orginal equations
y = 6(2) - 11
y = 12 - 11
y = 1
check.
-2(2) -3(1) = -7
-4 -3 = -7
-7 = -7
x = 2
y = 1
comment on my home page if this was helpful